OpenStudy (anonymous):

Sarah worked the following problem incorrectly. Part 1) In which line was the mistake made? Part 2) Show your work and give the correct answer? Line 1 40 ÷ 5(3 − 1) − 4 Line 2 40 ÷ 5(2) − 4 Line 3 40 ÷ 10 − 4 Line 4 4 − 4 Line 5 0

OpenStudy (anonymous):

Is It Line 2?

OpenStudy (zarkon):

what do you believe the mistake is?

OpenStudy (anonymous):

Error In PEMDAS

OpenStudy (zarkon):

sure ...but what exactly

OpenStudy (anonymous):

Multiplying?

OpenStudy (zarkon):

the part where that went from 5(2) to 10?

OpenStudy (anonymous):

Yes.

OpenStudy (zarkon):

yes...they should have divided first

OpenStudy (anonymous):

Yay! I Got It Right!

OpenStudy (zarkon):

40 ÷ 5(2)=8*2=16

OpenStudy (zarkon):

should't you say then that it is line 3?

OpenStudy (zarkon):

line 2 is equal to line 1 but line 3 is not equal to line 2

OpenStudy (anonymous):

Ok, So It's Line 3

OpenStudy (zarkon):

your reasoning is correct though

OpenStudy (dmezzullo):

so what is it?

OpenStudy (dmezzullo):

line3... whats wrong bout it?

OpenStudy (dmezzullo):

? i need to know what's wrong bout the problem to lol